OpenStudy (anonymous):

NEED HELP WILL MEDAL AND FAN. what are the first four terms of the following sequence: An=2^n/3! for n (equivalent or equal to) 1.?

jimthompson5910 (jim_thompson5910):

the sequence is \[\Large A_{n} = \frac{2^n}{3!}\] right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

re-write the 3! on the bottom, I would think.

OpenStudy (solomonzelman):

3! = 1 * 2 * 3

OpenStudy (anonymous):

yes @jim_thompson5910

OpenStudy (solomonzelman):

I was thinking to rew-rite it as: \[a_n=\frac{1}{6}2^n\]

jimthompson5910 (jim_thompson5910):

To find the first term \(\Large A_{1}\), replace n with 1 and evaluate \[\Large A_{n} = \frac{2^n}{3!}\] \[\Large A_{1} = \frac{2^1}{3!}\] \[\Large A_{1} = \frac{2^1}{3*2*1}\] \[\Large A_{1} = \frac{2}{6}\] \[\Large A_{1} = \frac{1}{3}\]

jimthompson5910 (jim_thompson5910):

and yes, \[\Large A_{n} = \frac{2^n}{3!} = \frac{1}{6}*2^n\]

jimthompson5910 (jim_thompson5910):

you will do the same for n = 2, n = 3, n = 4 to get terms 2 through 4

OpenStudy (anonymous):

so first term is 1/3, second is 2/3, third is 4/3, and fourth is 8/3?

OpenStudy (solomonzelman):

Or, \[a_n=\frac{1}{3}2^{n-1}\]

OpenStudy (solomonzelman):

using which ever version of a_n, what is the second term (when n=2) ?

jimthompson5910 (jim_thompson5910):

sar12389, that is correct

OpenStudy (solomonzelman):

didn't notice (s)he posted them . lol

OpenStudy (anonymous):

another question, Find The first four terms of the following sequence : An=2^n

OpenStudy (solomonzelman):

a(1) (when n=1) is?

jimthompson5910 (jim_thompson5910):

same idea, just different formula

OpenStudy (anonymous):

i tired it and i got 2,4,6, and 16

jimthompson5910 (jim_thompson5910):

2^3 is NOT equal to 6

jimthompson5910 (jim_thompson5910):

you mixed that up with 2*3

OpenStudy (anonymous):

i meant 8 instead of 6 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

2,4,8,16 is correct

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

np